### v3

# SimSort: A Data-Driven Framework for Spike Sorting by Large-Scale Electrophysiology Simulation

Abstract
Spike sorting is an essential process in neural recording, which identifies and separates electrical signals from individual neurons recorded by electrodes in the brain, enabling researchers to study how specific neurons communicate and process information. Although there exist a number of spike sorting methods which have contributed to significant neuroscientific breakthroughs, many are heuristically designed, making it challenging to verify their correctness due to the difficulty of obtaining ground truth labels from real-world neural recordings. In this work, we explore a data-driven, deep learning-based approach. We begin by creating a large-scale dataset through electrophysiology simulations using biologically realistic computational models. We then present SimSort, a pretraining framework for spike sorting. Trained solely on simulated data, SimSort demonstrates zero-shot generalizability to real-world spike sorting tasks, yielding consistent improvements over existing methods across multiple benchmarks. These results highlight the potential of simulation-driven pretraining to enhance the robustness and scalability of spike sorting in experimental neuroscience.
1 Introduction
Understanding the complex computations performed by the brain requires insight into the activity of individual neurons lewickiReviewMethodsSpike1998; buzsakiLargescaleRecordingNeuronal2004, which is crucial for exploring how information is encoded, processed, and transmitted within neural circuits, as well as decoding the brain’s functional dynamics quianquirogaExtractingInformationNeuronal2009; joshiDynamicSynchronizationHippocampal2023; sarkarAdvancedSpikeSorting2024. Recent advances in neural recording technologies have enabled capturing the activity of neurons across multiple regions of the brain with high spatial and temporal precision chungFullyAutomatedApproach2017; steinmetzNeuropixels20Miniaturized2021a; chungHighDensityLongLastingMultiregion2019; hongNovelElectrodeTechnologies2019. Notably, extracting meaningful information from recordings relies on a critical step known as spike sorting.
Spike sorting is the process of extracting and identifying neural activity from extracellular recordings. It involves two main steps (Fig. 1): spike detection, which extracts spike events from background noise, and spike identification, which assigns these detected spikes to individual neurons lefebvreRecentProgressMultielectrode2016. Spike sorting is indispensable for transforming raw electrical signals into interpretable data that reveal the firing patterns of individual neurons. Accurate spike sorting is essential for linking neural activity to behavior, understanding the functional organization of neural circuits sibilleHighdensityElectrodeRecordings2022, and uncovering mechanisms underlying various sensory and cognitive processes liuDecodingCognitionSpontaneous2022. Furthermore, its role extends to translational applications, such as improving neural prosthetics and developing closed-loop BCIs, where spike sorting is critical for achieving precise neural decoding and control park128ChannelFPGABasedRealTime2017; hanLiveDemonstrationEfficient2024.
For a long time, spike sorting predominantly relied on heuristic statistics and machine learning approaches lewickiReviewMethodsSpike1998; quirogaUnsupervisedSpikeDetection2004; rossantSpikeSortingLarge2016; pachitariuFastAccurateSpike2016; chungFullyAutomatedApproach2017; hilgenUnsupervisedSpikeSorting2017; pachitariuSpikeSortingKilosort42024. While these methods have promoted neuroscience research, they exhibit several key limitations. First, their sorting results are sensitive to parameter settings and post-processing, which depend on the experimenter’s expertise and must be customized for each dataset. Moreover, these approaches lack a data-driven foundation, limiting their scalability and adaptability across diverse experimental settings, particularly in low signal-to-noise ratio (SNR) and high-variability scenarios, where their performance declines significantly. Recently, deep learning-based methods, such as YASS leeYASSAnotherSpike2017 and CEED vishnubhotlaRobustGeneralizableRepresentations2023, have attempted to improve spike sorting by adopting data-driven approaches to enhance detection and clustering accuracy. However, their generalizability and practical use are hindered by the limited training datasets.
To address these limitations, we highlight the importance of using massive training data with ground-truth annotations to achieve more reliable and robust spike sorting. Furthermore, we aim to design a deep learning framework optimized for spike sorting and pretrain models for practical usage.
Therefore, in this work, we first generated a large-scale labeled dataset to address the scarcity of ground-truth data for spike sorting. Building on this, we present SimSort, a framework that utilizes data-driven approaches to enable fully automated spike sorting. By pretraining a spike detection model on the large-scale dataset, SimSort achieved notable improvements in detection accuracy and adaptability compared to commonly used threshold-based methods. Additionally, SimSort incorporated contrastive learning to enhance waveform feature representations, achieving better robustness against noise. We evaluated SimSort in zero-shot settings on publicly available datasets without fine-tuning, showing its effectiveness in spike sorting tasks compared to existing methods.
The key contributions of the SimSort framework include:
1) A publicly available, large-scale labeled dataset to help address the scarcity of ground-truth data, supporting the development and evaluation of learning-based spike sorting methods.
2) A pretraining paradigm for spike sorting with a large-scale simulation dataset. The results, for the first time, demonstrate successful zero-shot transfer from simulated to real-world spike sorting tasks.
3) An off-the-shelf pretrained model for neuroscientists to use***See https://SimSortTool.github.io for the model, code and usage instructions., enabling fully automated spike sorting on Tetrode recordings without dependence on hand-tuned parameters.
Figure 1: Pipeline of spike sorting, which consists of two main steps: spike detection and spike identification. For spike detection, typical spike sorting algorithms (left) utilize threshold-based detector relies on fixed voltage thresholds on each channel, and use non-learning method like performing PCA-based clustering on concatenated waveforms for spike identification. Those approaches are sensitive to noise and require manual parameter tuning. In contrast, our proposed framework SimSort (right) use a neural network-based detector replaces the threshold method for spike detection, enhancing robustness and generalization. For spike identification, the feature embeddings of multi-channel waveforms learned from contrastive learning to improve clustering accuracy.
2 Preliminaries
The goal of spike sorting (Figure 1) is to extract single-neuron spiking activity from extracellular recordings, where the activities of an unknown number of neurons are mixed together. Mathematically, spike sorting is a blind source separation problem buccinoSpikeSortingNew2022 and can be formally defined as follows: Let $\mathbf{V}\in\mathbb{R}^{T\times C}$ represent the extracellular recording (voltage) across $C$ electrode channels over $T$ time points. The goal is to infer a set of spike times $\{t_{k}\}_{k=1}^{K}$ and corresponding neuron labels $\{y_{k}\}_{k=1}^{K}$, where $y_{k}\in\{1,\dots,N\}$ represents the neuron identity, and $N$ is the total number of detected neurons.
3 Related Work
Early spike sorting pipelines typically used thresholds to detect spike events, followed by dimensionality reduction (e.g., PCA or $t$-SNE) and clustering FRS1901LIIIOL; maatenVisualizingDataUsing2008; veerabhadrappaCompatibilityEvaluationClustering2020. Although this pipeline was straightforward, it often suffered from noise susceptibility, inaccurate detection of low SNR events, and reliance on manual parameter tuning.
Deep learning approaches have been adopted to replace or enhance certain steps in traditional pipelines, aiming to improve robustness, automation, and adaptability to diverse experimental conditions. Autoencoders have been explored as a method for dimensionality reduction baldiAutoencodersUnsupervisedLearning2012; wuLearningSortFewshot2019; eomDeeplearnedSpikeRepresentations2021. YASS leeYASSAnotherSpike2017 employed a convolutional neural network for spike detection and waveform cleaning, thereby mitigating clustering errors caused by distorted waveforms. CEED vishnubhotlaRobustGeneralizableRepresentations2023 applied a contrastive learning framework to enforce invariances to amplitude fluctuations, noise, and channel subset changes in the extracted embeddings. However, these methods were limited by their reliance on restricted training datasets, potentially reducing their generalization across diverse experimental conditions.
In particular, YASS depended on high-quality prior training data and assumed a consistent experimental setup with validated sorting results, limiting its utility in scenarios where training data were scarce or recording conditions varied significantly.
Similarly, CEED had its limitations: (1) its invariance assumptions may not generalize to datasets with unaccounted variability; (2) it relied on KiloSort2.5-processed data, potentially inheriting inaccuracies from these analyses; (3) its training and testing datasets were narrowly scoped, originating from similar experimental conditions, which may constrain broader applicability; and (4) it did not optimize spike detection or provide a complete spike sorting pipeline.
To overcome these limitations, we propose a data-driven approach to develop a fully automated spike sorting pipeline that does not rely on manually defined parameters, demonstrating improved generalization across diverse datasets and recording conditions.
4 Methods
4.1 Dataset Creation
To address the scarcity of labeled spike data, we generated a large-scale synthetic dataset using biophysically detailed neuron models from the Blue Brain Project (BBP) hayModelsNeocorticalLayer2011; markramReconstructionSimulationNeocortical2015; laquitaineSpikeSortingBiases2025, covering 206 neuron models from layers 1–6 of the juvenile rat somatosensory cortex. These models capture diverse morphologies and electrophysiological dynamics across 30 neuronal types.
Electrophysiology Simulation
[…]
Figure 2: Overview of the large-scale simulated extracellular dataset generation process. (a) Multi-compartment neuron models simulate detailed neuronal morphologies and electrophysiological properties, incorporating realistic ion channel dynamics. (b) Noise generated by stochastic process is injected into the somatic compartment to induce stochastic intracellular action potential firing, and extracellular signals are recorded using virtual electrodes placed near the neurons in a simulated environment. (c) The resulting multi-channel extracellular recordings capture diverse and realistic neural activity.
Dataset Preparation
We generated a large-scale dataset consisting of 8192 simulated recording trials, representing continuous neuronal activities from more than 40,000 individual neurons. Intracellular spike timestamps were recorded as ground-truth for extracellular spikes, providing precise annotations for model training and evaluation.
The dataset was structured into two primary subsets:
Continuous Signal Dataset: This subset comprised full-length extracellular recordings across all electrode channels. It was designed for training spike detection models and evaluating the overall performance of spike sorting pipelines.
Spike Waveform Dataset: This subset contained spike waveforms extracted from each ground-truth units. It was used for training and evaluating the spike identification model.
We used signals simulated from BBP layers 1–5 neuron models for training and validation, reserving signals from BBP layer 6 neuron models exclusively for evaluation, ensured that the test set comprises previously unseen neuron types and configurations, allowing us to effectively assess the model’s generalization capabilities.
4.2 Spike Detection
We formulate the spike detection task as identifying the temporal segments $\{t_{k}\}$ corresponding to putative neural spikes from the raw input $\mathbf{V}\in\mathbb{R}^{T\times C}$, where $T$ represents the number of time points and $C$ denotes the number of electrode channels. The raw signal V first underwent two preprocessing steps: bandpass filtering and spatial whitening. Then, the processed signal was subsequently provided as input to the spike detection model.
Figure 3: Spike detection model in SimSort.
We utilized a Transformer-based architecture vaswaniAttentionAllYou2017 for spike detection, trained on the simulated continuous signal dataset detailed in Sec. 4.1. For robustness and generalization, data augmentation was integrated into the training process, referring to Fig. 3. We employed a binary cross-entropy loss for the spike detection model:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{T}\sum_{t=1}^{T}\Big(w_{p}\cdot y_{t}\log\hat{y}_{t}+(1-y_{t})\log(1-\hat{y}_{t})\Big),$ (1)
where the binary label $y_{t}$ denoted spike events, $\hat{y}_{t}$ was the predicted probability, $T$ was trial length, and $w_{p}$ was a hyperparameter to reweigh the positive predictions given the general scarcity of spikes.
Data Augmentation: To improve robustness and generalization in spike detection, data augmentation was applied to each input signal segment $V^{\prime}\in\mathbb{R}^{T\times C}$ (Fig. 3). A subset of channels $C^{\prime}\subseteq\{1,2,\dots,C\}$ was randomly selected with a probability $p$ to apply augmentation. Various augmentations, such as adding noise, amplitude scaling, and temporal jitter, were applied to enhance the diversity of training data. The detailed implementation of these augmentations is provided in Appendix A.
4.3 Spike Identification
We formulate spike identification as embedding each identified spike waveform $X_{i}\in\mathbb{R}^{L\times C}$, the voltage trace (waveform) of spike $i$, into a latent space to learn robust representations using contrastive learning. Then the learned embeddings are clustered into several groups, indicating putative neurons.
The spike identification model was trained on the simulated spike waveform dataset (Sec. 4.1). The model consisted of several key components, as illustrated in Fig. 4 and described as follows.
Figure 4: Spike identification model in SimSort.
Contrastive Learning: We leveraged contrastive learning to learn representations that capture the relationships between spike waveforms (Fig. 4). Contrastive learning sought to map neural data into an embedding space where related examples were positioned nearby, while unrelated examples were placed further apart dorkenwaldMultilayeredMapsNeuropil2023; yuVivoCelltypeBrain2024. For each waveform $X_{i}$, we constructed a triplet by selecting $X_{i}^{+}$ with the same label and $X_{i}^{-}$ with a different label. The objective was schroffFaceNetUnifiedEmbedding2015:
$\mathcal{L}_{\text{triplet}}=\max\big(0,\|f(X_{i})-f(X_{i}^{+})\|_{2}^{2}-\|f(X_{i})-f(X_{i}^{-})\|_{2}^{2}+\alpha\big),$ (2)
where $f(X_{i})$ denoted the embedding of $X_{i}$, and $\alpha$ was a margin enforcing separation between positive-negative pairs. Additionally, we uniformly applied data augmentation to each triplet waveforms. The augmentation methods followed those in Sec. 3, with adjusted parameters (Appendix A) for better waveform representation, ensuring robustness to noise, amplitude variations, and temporal jitter.
Denoising: The augmented waveform $X_{i}^{\prime}\in\mathbb{R}^{L\times C}$ then underwent a denoising step using Singular Value Decomposition (SVD) eckartApproximationOneMatrix1936; cunninghamDimensionalityReductionLargescale2014. The waveform was first reshaped and decomposed as $X_{i}^{\prime}=U\Sigma V^{\top}$, where $U$, $\Sigma$, and $V$ represented the singular vectors and singular values. The first $k$ components were retained to reconstruct the waveform.
Encoder: The denoised waveform $X_{i,\text{recon}}^{\prime}$ was fed into a Gated Recurrent Unit (GRU) encoder choLearningPhraseRepresentations2014 (Fig. 4), which processed the input sequentially. The final hidden state $h_{T}$ served as the learned representation $f(X_{i})$, encapsulating the spatiotemporal features of the waveform for subsequent clustering tasks.
Clustering: After obtaining the waveform representation, we applied dimensionality reduction using UMAP mcinnesUMAPUniformManifold2020 before clustering (Fig. 4). We tested two clustering approaches: a parametric algorithm, Gaussian Mixture Model (GMM) dempsterMaximumLikelihoodIncomplete1977, and a non-parametric one, Mean Shift (MS) chengMeanShiftMode1995.
4.4 Overall Spike Sorting Pipeline
We established a cascaded spike sorting pipeline (Fig. 1). During inference, raw multi-channel signals were processed through spike detection, waveform extraction, embedding generation, and clustering, yielding distinct neuronal units and their spike timestamps.
5 Results
With the large-scale dataset and a pretraining framework, we try to answer the following research questions (RQ).
RQ1: Can the model trained on our simulated dataset generalize to real-world spike sorting tasks?
RQ2: If the model performs well, what are the crucial underlying components?
RQ3: Does data scale hold the key of effective spike sorting?
5.1 Benchmarks
Since the models of SimSort has been trained on our simulated dataset, readers must be curious about how well it can generalize to real-world spike sorting tasks. To empirically investigate this problem, we respectively evaluate SimSort in the following benchmarks:
  • Simulated test dataset (with ground truth): BBP L6 dataset from our simulation (using different neuronal models from the training dataset).
  • Real-world data-based synthetic datasets (with label): Hybrid (maglandSpikeForestReproducibleWebfacing2020), WaveClus (martinezRealisticSimulationExtracellular2009), IBL Neuropixels (vishnubhotlaRobustGeneralizableRepresentations2023). We evaluate both zero-shot (using only our simulated dataset) and fine-tuned (using also the dataset) performance of SimSort.
  • Real-world neural recording (w/o ground truth, indirect validation): extracellular recordings collected from the mouse primary visual cortex during drifting grating stimulation.
The detailed information of benchmark datasets, hyper-parameters of models, and formulation of the evaluation metrics can be found in Appendix A.
Evaluation Metrics. To evaluate the performance of the overall spike sorting and spike detection tasks, we used Accuracy, Recall, and Precision as the evaluation metrics. Note that these metrics are specifically defined for the spike sorting context, where they measure the alignment between detected spikes and ground truth spike events, rather than their conventional definitions in classification tasks. For the spike identification task, we used the Adjusted Rand Index (ARI) hubert1985comparing to measure clustering accuracy relative to the ground truth.
Table 1: Spike sorting results on BBP L6 dataset. The results for other methods were obtained through SpikeInterface buccinoSpikeInterfaceUnifiedFramework2020. Values are as mean $\pm$ S.E.M.
Methods BBP L6 (20 trials)
Accuracy Recall Precision
KiloSort 0.49 $\pm$ 0.05 0.51 $\pm$ 0.05 0.53 $\pm$ 0.05
KiloSort2 0.51 $\pm$ 0.06 0.53 $\pm$ 0.06 0.53 $\pm$ 0.06
MountainSort4 0.84 $\pm$ 0.03 0.84 $\pm$ 0.03 0.93 $\pm$ 0.03
MountainSort5 0.66 $\pm$ 0.06 0.68 $\pm$ 0.06 0.79 $\pm$ 0.07
[…]
5.2 Spike Sorting Results on simulated dataset
We first examined Simsort on a simulated test set (BBP L6). In Table 1, we evaluated SimSort alongside widely used algorithms, including KiloSort pachitariuFastAccurateSpike2016, KiloSort2, MountainSort4 chungFullyAutomatedApproach2017, and MountainSort5, implemented via the SpikeInterface pipeline buccinoSpikeInterfaceUnifiedFramework2020, on the simulated BBP L6 dataset. SimSort achieved the best performance across all metrics, demonstrating its robustness and strong zero-shot generalization ability under idealized, simulated conditions.
Table 2: Spike sorting results on Hybrid dataset. Data for other methods were sourced from SpikeForest maglandSpikeForestReproducibleWebfacing2020. Values are mean $\pm$ S.E.M. Best results are in bold. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
Methods Hybrid-static (SNR>>3, 9 recordings) Hybrid-drift (SNR>>3, 9 recordings)
Accuracy Recall Precision Accuracy Recall Precision
HerdingSpikes2 0.35 $\pm$ 0.01 0.44 $\pm$ 0.02 0.53 $\pm$ 0.01 0.29 $\pm$ 0.01 0.37 $\pm$ 0.02 0.48 $\pm$ 0.02
IronClust 0.57 $\pm$ 0.04 0.81 $\pm$ 0.01 0.60 $\pm$ 0.04 0.54 $\pm$ 0.03 0.71 $\pm$ 0.02 0.65 $\pm$ 0.03
JRClust 0.47 $\pm$ 0.04 0.63 $\pm$ 0.02 0.59 $\pm$ 0.03 0.35 $\pm$ 0.03 0.48 $\pm$ 0.03 0.57 $\pm$ 0.02
KiloSort 0.60 $\pm$ 0.02 0.65 $\pm$ 0.02 0.72 $\pm$ 0.02 0.51 $\pm$ 0.02 0.62 $\pm$ 0.01 0.72 $\pm$ 0.03
KiloSort2 0.39 $\pm$ 0.03 0.37 $\pm$ 0.03 0.51 $\pm$ 0.03 0.30 $\pm$ 0.02 0.31 $\pm$ 0.02 0.57 $\pm$ 0.04
MountainSort4 0.59 $\pm$ 0.02 0.73 $\pm$ 0.01 0.73 $\pm$ 0.03 0.36 $\pm$ 0.02 0.57 $\pm$ 0.02 0.61 $\pm$ 0.03
SpykingCircus 0.57 $\pm$ 0.01 0.63 $\pm$ 0.01 0.75 $\pm$ 0.03 0.48 $\pm$ 0.02 0.55 $\pm$ 0.02 0.68 $\pm$ 0.03
Tridesclous 0.54 $\pm$ 0.03 0.66 $\pm$ 0.02 0.59 $\pm$ 0.04 0.37 $\pm$ 0.02 0.52 $\pm$ 0.03 0.55 $\pm$ 0.04
[…]
Table 3: Spike sorting results on WaveClus dataset. The results for other methods were obtained through SpikeInterface buccinoSpikeInterfaceUnifiedFramework2020. Values are mean $\pm$ S.E.M. Note that SimSort did not use this dataset for training, indicating a zero-shot generalization.
Methods Easy (12 recordings) Difficult (8 recordings)
Accuracy Recall Precision Accuracy Recall Precision
KiloSort 0.54 $\pm$ 0.12 0.46 $\pm$ 0.15 0.62 $\pm$ 0.13 0.17 $\pm$ 0.07 0.23 $\pm$ 0.10 0.18 $\pm$ 0.07
KiloSort2 0.61 $\pm$ 0.09 0.66 $\pm$ 0.09 0.67 $\pm$ 0.09 0.10 $\pm$ 0.07 0.12 $\pm$ 0.09 0.10 $\pm$ 0.07
MountainSort4 0.73 $\pm$ 0.08 0.79 $\pm$ 0.07 0.79 $\pm$ 0.08 0.48 $\pm$ 0.16 0.49 $\pm$ 0.16 0.52 $\pm$ 0.17
MountainSort5 0.71 $\pm$ 0.09 0.74 $\pm$ 0.08 0.79 $\pm$ 0.10 0.42 $\pm$ 0.15 0.46 $\pm$ 0.15 0.46 $\pm$ 0.16
[…]
5.3 Spiking Sorting results on real-world data-based synthetic datasets
zero-shot generalization
Given SimSort’s strong performance on simulated data, we extended our evaluation to real-world data-based datasets to further test its zero-shot generalizability and robustness (RQ1). On the Hybrid dataset, we compared SimSort against several algorithms via SpikeForest, including HerdingSpikes2 hilgenUnsupervisedSpikeSorting2017, IronClust, JRClust, KiloSort pachitariuFastAccurateSpike2016, KiloSort2, MountainSort4 chungFullyAutomatedApproach2017, SpykingCircus, and Tridesclous. The results are summarized in Table 2 and Fig. S1. On this dataset, SimSort demonstrated the highest Accuracy and Precision in static recordings and achieved the best Accuracy and Recall under drift conditions, although it slightly trailed KiloSort in Precision.
For WaveClus dataset, we used SpikeInterface pipeline to run KiloSort, KiloSort2, MountainSort4, and MountainSort5. Results are provided in Table 3 and Fig. S1. SimSort performed reliably across both easy and difficult subsets, showing a pronounced advantage on the challenging “difficult” set. This performance was likely due to its ability to handle the unique challenges of this subset, such as lower SNRs and reduced inter-class waveform variability, which resulted in significant overlaps between units. These results underscored SimSort’s effectiveness in noisy and complex scenarios.
Fine-tuning
Beyond demonstrating SimSort’s zero-shot transfer ability on unseen datasets, we also investigate its performance after fine-tuning on a small amount of data from the target domain (Table 4). While the overall improvements in sorting metrics are modest, we observe notable changes in spike detection performance, particularly in the balance between Precision and Recall. These changes suggest that fine-tuning helps the model adapt to dataset-specific waveform and noise characteristics, enhancing its practicality with minimal labeled data.
Table 4: Performance metrics of SimSort after fine-tuning on different datasets.The hybrid-static and hybrid-drift datasets were fine-tuned on recordings 1-3 and tested on recordings 4-9 (train/test set split). The waveclus -easy and -difficult datasets were fine-tuned on subsets easy1/diff1 and tested on easy2/diff2, respectively. The IBL Neuropixels dataset was fine-tuned on 100 neurons and evaluated on 50 test sets, each with 10 randomly selected neurons.
Dataset Model Spike Detection (Acc/Rec/Prec) Spike Identification (ARI) Spike Sorting (Acc/Rec/Prec)
Hybrid-static SimSort 0.71±0.02 / 0.83±0.02 / 0.83±0.02 0.90±0.02 0.59±0.03 / 0.65±0.03 / 0.76±0.02
SimSort-FT 0.75±0.03 / 0.89±0.02 / 0.82±0.02 0.92±0.02 0.58±0.03 / 0.65±0.03 / 0.73±0.02
Hybrid-drift SimSort 0.67±0.02 / 0.80±0.02 / 0.81±0.01 0.88±0.03 0.54±0.09 / 0.61±0.09 / 0.67±0.07
SimSort-FT 0.73±0.03 / 0.85±0.03 / 0.83±0.01 0.88±0.02 0.56±0.04 / 0.63±0.03 / 0.69±0.03
Waveclus-easy SimSort 0.59±0.02 / 0.99±0.00 / 0.59±0.02 0.91±0.01 0.70±0.11 / 0.79±0.07 / 0.84±0.09
SimSor-FT 0.94±0.01 / 0.96±0.00 / 0.98±0.01 0.91±0.01 0.73±0.08 / 0.79±0.05 / 0.91±0.04
Waveclus-difficult SimSort 0.66±0.08 / 0.99±0.00 / 0.67±0.08 0.88±0.01 0.76±0.14 / 0.84±0.09 / 0.84±0.12
SimSort-FT 0.95±0.01 / 0.96±0.00 / 0.99±0.01 0.87±0.03 0.77±0.14 / 0.82±0.10 / 0.86±0.13
IBL Neuropixels SimSort N/A 0.49±0.09 N/A
SimSort-FT N/A 0.53±0.09 N/A
5.4 Real-World Neural Recording
Beyond quantitative benchmarks, we further assessed SimSort in a real experimental setting. Specifically, we applied it to extracellular recordings (collected in our wet lab, see Appendix A.1.2) from the mouse primary visual cortex obtained using tetrodes during presentation of full-field drifting gratings in 8 orientations (Fig. 5a). From a representative recording session, SimSort identified six putative single units (Fig. 5b). These units exhibited distinct spike waveforms and their autocorrelograms showed clear refractory period structure(Fig. 5c), suggesting good isolation. We further analyzed the orientation tuning of these units. Several exhibited selective responses, with global orientation selectivity index (gOSI) values exceeding 0.2 (Fig. 5d; Appendix A for definition). Across 10 tetrode recordings, SimSort isolated a total of 40 units, with a mean gOSI of 0.27 ± 0.14 (mean ± S.D., Fig. S17). These results demonstrate the practical applicability of SimSort to real neural recordings.
Figure 5: Evaluation of SimSort on real experimental data (See Fig. S17 for additional recording results). (a) Experimental setup: extracellular recordings from the mouse primary visual cortex during drifting gratings in eight directions. (b) Visualization of SimSort’s sorting results. (c) Autocorrelograms and average waveform of each unit. (d) Orientation tuning curves and global orientation selectivity index (gOSI) for each unit.
5.5 Detailed Analysis
To explain why our framework achieved higher overall spike sorting accuracy (RQ2), we dissected SimSort and analyzed each component, i.e., spike detection and spike identification models. Furthermore, we presented some case studies of the latent embedding of spikes from multiple neurons.
Spike Detection
[…]
Table 5: Spike detection and identification results on the Hybrid dataset. Values are mean $\pm$ S.E.M.
(a) Spike detection results. The threshold detector used the best-performing threshold by grid search.
Dataset Method Accuracy Recall Precision
[…]
(b) Spike identification results. All results averaged across 20 repeats of GMM.
Dataset PCA $t$-SNE UMAP Ours
Hybrid-static 0.79 $\pm$ 0.04 0.88 $\pm$ 0.02 0.88 $\pm$ 0.02 0.91 $\pm$ 0.02
Hybrid-drift 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02 0.85 $\pm$ 0.01 0.89 $\pm$ 0.03
Spike Identification
After detecting spikes, our spike identification model extracted waveform representations in the latent space, enabling subsequent clustering. To effectively demonstrate the superiority of our identification model, we assembled a waveform dataset by gathering waveform data from each neuron in every hybrid recording. Using this dataset, we compared the 2D UMAP embedding of our model’s inferred representations with features extracted directly from waveforms using widely adopted dimensionality reduction techniques, including PCA FRS1901LIIIOL, $\boldsymbol{t}$-SNE maatenVisualizingDataUsing2008, and UMAP mcinnesUMAPUniformManifold2020. For evaluation, we applied a parametric clustering method, GMM dempsterMaximumLikelihoodIncomplete1977, to compute the Adjusted Rand Index (ARI), providing a quantitative measure of the reliability of the representations.
As shown in Table 5(b), SimSort substantially outperformed these conventional methods under both static and drift conditions, achieving the highest average ARI values of 0.91 and 0.89, respectively. These results revealed the robustness and effectiveness of our approach in deriving meaningful and reliable representations from extracellular data.
Additionally, we compared SimSort with CEED vishnubhotlaRobustGeneralizableRepresentations2023, a recent framework for representation learning on extracellular data, using the author-provided model. Note that CEED is trained on datasets derived from IBL DanLab DY016 and DY009 Neuropixels recordings, which are more closely related to our test sets. We evaluated both SimSort and CEED on fifty test sets from the IBL Neuropixels dataset (specifically derived from the IBL CortexLab KS046 recording). As presented in Table S4, Fig. S2S3, our method, without training on the dataset, achieved performance comparable to CEED.
Case Studies: In Fig. 6, we visualized the results of spike identification on an example recording in the Hybrid static subset. SimSort’s embeddings (first column) demonstrated superior clustering performance, with a GMM clustering ARI score of 0.90, closely aligning with true labels. In contrast, UMAP, $t$-SNE, and PCA embeddings (second to fourth columns) yielded lower clustering scores of 0.80, 0.78, and 0.67, respectively. SimSort embeddings showed clear separation between neuron clusters, while other methods exhibited significant overlap, highlighting the advantages of SimSort in learning robust and discriminative features.
Impact of Dimensionality Reduction and Clustering Methods: To evaluate the adaptability of SimSort, we conducted a study to assess the impact of different dimensionality reduction techniques (PCA, $t$-SNE, UMAP, and None) and clustering algorithms (GMM, KMeans macqueenMethodsClassificationAnalysis1967, Mean Shift chengMeanShiftMode1995, and Spectral Clustering ngSpectralClusteringAnalysis2001). Dimensionality reduction was applied to latent features extracted from waveforms, with an additional setup where clustering was performed directly on the raw latent features without dimensionality reduction (None). As illustrated in Fig. S4, the combination of UMAP and Mean Shift achieved the highest accuracy on both hybrid static and drift subsets. Across many other combinations, SimSort also maintained competitive performance, demonstrating its flexibility and robustness to various configurations.
Figure 6: Visualized results of spike identification on an example recording in the Hybrid static subset. See Appendix C for visualization results for additional samples.
5.6 Scaling Law of Data Size
Figure 7: Performance scaling with increasing training data size on the Hybrid-drift and -static.
As we emphasized the importance of a large-scale dataset throughout this work, it was interesting to evaluate how SimSort’s performance scaled with the size of the training dataset (RQ3). Specifically, we evaluated performance across sorting accuracy, detection accuracy, and identification ARI score, capturing how each metric varied with the growing size of the training dataset, which ranged from $2^{8}$ to $2^{13}$ trials. As Fig. 7 shows, larger datasets led to almost consistent performance improvements on both hybrid static and drift subsets, reflecting that the large-scale dataset holds the key of SimSort’s effectiveness.
6 Conclusion
In this paper, we presented SimSort, a data-driven framework for automated spike sorting, built upon a biologically realistic large-scale extracellular dataset. SimSort achieves competitive or superior performance compared to commonly used spike sorting algorithms on both simulated and real-world benchmarks, demonstrating its robustness and scalability. Evaluations further highlighted its zero-shot transfer capability.
SimSort has limitations. The models were trained on 4-channel tetrode data. While SimSort shows promising performance on tasks with high-density probes (IBL Neuropixels), future work should extend the training with diverse electrode geometries.
Acknowledgments and Disclosure of Funding
This work was supported by AI for Science Foundation of Fudan University (FudanX24AI046 to Y.G.) and Microsoft Research.
Broader Impacts
Our work aims to advance the methodology of data processing in neuroscience experiments. It does not directly lead to any negative societal impact.
Appendix A Implementation Details of SimSort
A.1 Data Preparation
A.1.1 Train datasets
In Sec. 4.1, we introduced the preparation of our simulated training dataset, divided into the continuous signal dataset and the spike waveform dataset.
The continuous signal dataset was constructed by concatenating multiple simulated trials (e.g., 8192 trials), each containing 6,000,000 timesteps (dt = 0.1 ms) of extracellular data. Spike labels were derived by identifying peak times of intracellular action potentials and assigning binary labels to the corresponding region. This dataset, therefore, provided labeled, continuous extracellular signals for training the detection model and evaluating the overall spike sorting performance of SimSort.
The spike waveform dataset was constructed by extracting waveform segments centered on labeled spike events. For each spike, a 60-timestep waveform was segmented from the extracellular signals, and up to 400 such waveforms were collected per unit. Each waveform was then paired with its corresponding unit label, creating a dataset dedicated to training and evaluating the identification model.
A.1.2 Test datasets
BBP L6 dataset (simulation). This dataset comprised 20 simulated trials, generated as stated in Sec. 4.1, using BBP layer 6 neuron models. It provided 5 ground-truth units per recording for evaluating spike sorting performance.
Hybrid dataset. This dataset was composed of recordings provided by Spikeforest [maglandSpikeForestReproducibleWebfacing2020]. These recordings were generated by Kilosort, with waveform templates recorded at a 5 µm electrode spacing and  1/f background noise. It consisted of two subsets: 9 static tetrode recordings and 9 drift tetrode recordings that simulated sinusoidal probe movements. For each recording, we selected units with an SNR greater than 3 (unit IDs meeting this criterion were provided by SpikeForest to ensure a fair comparison), comprising 10–15 ground-truth units per recording. The dataset we prepared supported two modes of use. To evaluate overall spike sorting and detection performance, we used the raw data as input. For assessing identification performance, we constructed a waveform dataset by extracting 500 waveforms (60-timesteps length) per unit.
WaveClus dataset. This dataset consisted of four subsets (8 recordings in Easy1, 4 recordings in Easy2, 4 recordings in Difficult1, and 4 recordings in Difficult2). Each recording was generated by combining spike waveform templates derived from experimental recordings with background noise, with noise levels ranging from 0.05 to 0.4, and contained 3 ground-truth units. To increase the overall difficulty and adapt the original single-channel recording for tetrode analysis, we preprocessed the signals by duplicating them and modifying them with small random noise (standard deviation 0.01), amplitude scaling (ranging from 1 to 0.5), and random temporal shifts (-5 to +5 samples).
IBL Neuropixels dataset. This dataset was derived from a real-world extracellular recording, CortexLab KS046, published by the International Brain Laboratory (IBL). This recording was captured with a Neuropixels 1.0 probe across multiple brain regions from a mouse performing a decision-making behavior task. To compare spike identification performance with CEED, we prepared this dataset following the method described in the CEED paper (CEED trained and evaluated datasets derived from DanLab DY016 and DY009 Neuropixels recordings in the IBL database). We extracted waveforms from 400 units detected by Kilosort 2.5 and created fifty test sets for comparison between SimSort and CEED. Each test set contained 10 random neurons and 100 spikes per neuron. Note that CEED inputs 11-channel data into the model, while SimSort processes data from the center 4 channels.
Real-world Neural Recording. These recordings were collected by our wet lab (approved by research ethics committee). A 16-channel tetrode was implanted unilaterally in the primary visual cortex (V1) of mice. During the recording sessions, the mice were head-fixed, and the contralateral eye was stimulated with drifting grating stimuli of 8 different orientations. The stimuli were presented in a randomized sequence over 8 repeated trials. Each grating had a spatial frequency of 0.04 cycles per degree (cpd) and a temporal frequency of 2 Hz, lasting for 1.5 seconds per trial.
To quantify the orientation tuning of individual neurons, we computed the global orientation selectivity index (gOSI), as defined in the equation below:
$gOSI=\frac{\left\|\sum R(\theta)e^{2i\theta}\right\|}{\sum R(\theta)}$
where $\theta$ is the direction of the moving grating and $R(\theta)$ is the neuronal response to direction $\theta$. The imaginary unit is denoted as $i$ whose square is 1-1. The gOSI ranges from 0 to 1, with 0 indicating no orientation selectivity (equal response to all directions), and 1 indicating maximal selectivity (response only to a single orientation).
Table S1: Details information of test datasets used for evaluation.
Datasets Num. Recordings Sample Rate (Hz) Num. Channels Duration (sec per rec.) Num. True Units (per rec.) Ground-Truth Determination
BBP L6 dataset 20 10000 4 600 5 Simulation
Hybrid dataset 18 30000 4 600 / 1200 10-15 Real waveforms + background noise
WaveClus dataset 20 24000 4 60 3 Real waveforms + background noise
IBL Neuropixels dataset / 30000 11 / 10 Sorted with KiloSort 2.5
Real-world Neural recording 10 40000 4 $\approx 250$ / No Ground-truth
A.2 Technical Details of Simulation
Intracellular Simulation. Intracellular electrophysiological activity was simulated using the NEURON package [hinesNEURONSimulationEnvironment1997, hinesNeuronToolNeuroscientists2001]. Pink noise, scaled to the rheobase (the minimum current needed to elicit an action potential in a neuron), was injected into each neuron’s soma to evoke stochastic action potentials, simulating biologically realistic background noise. The temporal resolution was set to 0.1 ms, and simulations spanned a total duration of 600 seconds. Neurons were initialized with a resting membrane potential of -70 mV and maintained at a physiological temperature of 34°C. Synaptic mechanisms and biophysical properties were dynamically compiled and loaded to ensure compatibility with the NEURON simulation environment.
For each neuron, a multi-compartmental model [hayModelsNeocorticalLayer2011, markramReconstructionSimulationNeocortical2015, gouwensSystematicGenerationBiophysically2018, billehSystematicIntegrationStructural2020] was used to calculate transmembrane currents [lindenLFPyToolBiophysical2014, hagenViSAPyPythonTool2015, hagenMultimodalModelingNeural2018, buccinoMEArecFastCustomizable2021]. The neuron was divided into compartments, each described as an equivalent electrical circuit. The dynamics of the membrane potential in each compartment $n$ were governed by Kirchhoff’s current law [holmesPassiveCableModeling2009], where the sum of all currents entering or leaving a node must equal zero. Considering that extracellular potential changed much slower than ion channel dynamics, we assumed it constant. The temporal evolution of the membrane potential $V_{n}$ in compartment $n$ was given by $g_{n,n+1}(V_{n+1}-V_{n})-g_{n-1,n}(V_{n}-V_{n-1})=C_{n}\frac{dV_{n}}{dt}+\sum_{j}I_{n}^{(j)},$ where $g_{n,n+1}$ and $g_{n-1,n}$ represented the conductances between neighboring compartments, $C_{n}$ was the membrane capacitance, and $\sum_{j}I_{n}^{(j)}$ accounted for ionic currents and any externally applied currents.
These transmembrane currents served as the source for extracellular potential modeling in the next simulation stage.
[…]
The extracellular potential at a given electrode site $\mathbf{r}_{e}$ was computed as the sum of contributions from all neuronal segments within the simulation space: $\phi(\mathbf{r}_{e},t)=\frac{1}{4\pi\sigma}\sum_{n=1}^{N}\sum_{m=1}^{M_{n}}\frac{I_{n,m}(t)}{|\mathbf{r}_{e}-\mathbf{r}_{n,m}|},$ where $I_{n,m}(t)$ was the transmembrane current of the $m$-th compartment of the $n$-th neuron, $\mathbf{r}_{n,m}$ was its position, $\sigma$ is the extracellular medium’s conductivity, and $|\mathbf{r}_{e}-\mathbf{r}_{n,m}|$ was the Euclidean distance between the segment and the electrode. Here, $N$ was the total number of neurons, and $M_{n}$ was the number of compartments in the $n$-th neuron. A reference electrode far from the source was assumed, setting $\phi=0$ at infinity.
A.3 Spike Detection
Data Augmentation
For each input segment $X_{i}\in\mathbb{R}^{T\times C}$, where $T$ represents the number of time points and $C$ denotes the number of electrode channels, a subset of channels $\mathcal{C}\subseteq\{1,2,\dots,C\}$ was randomly selected. A probability $p\in[0,1]$ determined whether the following augmentations were independently applied to each channel $k\in\mathcal{C}$:
$X^{\prime}_{i,k}=\begin{cases}X_{i,k}+\epsilon,&\text{AddWithNoise},\\ \beta\cdot X_{i,k},&\text{RandomAmplitudeScaling},\\ X_{i,k,\lfloor t\cdot n+\epsilon_{t}\rfloor},&\text{RandomTimeJitter}.\end{cases}$ (3)
where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise, $\beta\sim\mathcal{U}(1-s,1+s)$ is a random amplitude scaling factor, and $\epsilon_{t}\sim\mathcal{U}(-\delta,\delta)$ introduces small temporal shifts. The parameter $n$ is an up-sampling factor enabling finer time resolution.
Model-Based Detector
In Sec. 4.2, we introduced a model-based detector to predict spike events. These predictions were processed through the following steps:
1. Detection with Model: The model took as input a multi-channel signal $V\in\mathbb{R}^{t\times c}$, where $t$ was the number of time samples and $c$ was the number of channels. This input was processed simultaneously across all channels by the detection model, which predicted a sequence of binary labels $\hat{y}_{i}(t)$ for each time point:
$\hat{y}_{i}(t)=f_{\text{model}}(V),$ (4)
where $f_{\text{model}}$ represented the detection model. The output $\hat{y}_{i}(t)$ indicated the presence or absence of spike events at each time point.
2. Adjacent Peak Merging: The predicted labels $\hat{y}_{i}(t)$ were refined by merging adjacent peaks to reduce false positives and consolidate overlapping detections. To identify the most relevant peaks, the channel with the maximum summed absolute signal across all time points was selected:
$c^{*}=\arg\max_{c\in\{1,\dots,C\}}\sum_{t=1}^{T}|V(t,c)|,$ (5)
where $V(t,c)$ denoted the signal value at time $t$ for channel $c$. Using this peak channel, adjacent peaks were grouped within a window size $\Delta t$, and only the time point with the maximum absolute signal was retained within each group:
$t_{\text{max}}=\arg\max_{t\in[t_{s},t_{e}]}|V(t,c^{*})|,$ (6)
where $t_{s},t_{e}]$ denoted the boundaries of a group of adjacent peaks.
3. Spike Peak Extraction: After merging, the final spike peak positions $\{t_{k}\}$ were determined as the set of time points where $\hat{y}^{\prime}_{i}(t)=1$:
$t_{k}=\{t\,|\,\hat{y}^{\prime}_{i}(t)=1\}.$ (7)
Threshold-Based Detector
In Sec. 5.5, we compared our model-based detector with a threshold-based detector. To ensure a fair comparison, the processing steps for the threshold-based detector were designed to align closely with those of the model-based detector. The threshold-based method processed the signal $V\in\mathbb{R}^{t\times c}$, where $t$ was the number of time samples and $c$ was the number of channels. Each channel was processed independently, as follows:
1. Local Maxima Detection: For each channel $c$, the signal $V(t,c)$ was scanned to identify local maxima that satisfied the following conditions:
$V(t,c)>Th_{\text{single\_ch}},\quad V(t,c)>V(t-r,c),\quad V(t,c)>V(t+r,c),$ (8)
where $Th_{\text{single\_ch}}$ was the amplitude threshold, and $r\in\{1,\dots,\text{loc\_range}\}$. These conditions ensured that the detected peaks were above the threshold and were local maxima within the range $r$. Peaks identified by this criterion were further validated within a broader range long_range to ensure they represented significant spike events.
2. Adjacent Peak Merging: For each channel $c$, adjacent peaks were grouped within a window size $\Delta t$, and only the time point with the maximum absolute signal was retained within each group:
$t_{\text{max}}=\arg\max_{t\in[t_{s},t_{e}]}|V(t,c)|,$ (9)
where $t_{s},t_{e}]$ defined the boundaries of a cluster of adjacent peaks. This step ensured that closely spaced peaks were consolidated into a single spike.
3. Spike Peak Extraction: After processing all channels independently, the final detected spike positions for channel $c$ were determined as:
$t_{k}=\{t\,|\,V(t,c)=t_{\text{max}}\}.$ (10)
A.4 Spike Identification
Waveform Snippet Extraction. For each detected spike at position $t_{k}$, a waveform snippet $W_{k}\in\mathbb{R}^{L\times c}$ was extracted from the multi-channel continuous signal $V(t,c)$, where $L$ was the snippet length (number of time samples) and $c$ was the number of channels. The snippet was centered around $t_{k}$ and spanned a time window determined by the pre- and post-padding values:
$W_{k}=V(t_{k}-p_{\text{pre}}:t_{k}+p_{\text{post}},:),$ (11)
where $V(t,c)$ represented the input signal value at time $t$ for channel $c$, and $p_{\text{pre}}$ and $p_{\text{post}}$ were the padding values (in time samples) before and after the spike time $t_{k}$, respectively.
A.5 Definitions of Metrics
To evaluate the performance of spike sorting algorithms, we adopted standard metrics as defined in the context of spike sorting and ground-truth comparisons [maglandSpikeForestReproducibleWebfacing2020, buccinoSpikeInterfaceUnifiedFramework2020]. These metrics included accuracy, precision, and recall, which were derived based on the matching between the sorted spike train and the ground-truth spike train.
1. Matching Spike Events: For a given sorted spike train $k$ and ground-truth spike train $l$, the first step was to calculate the number of matching events ($n^{\text{match}}_{l,k}$). A sorted spike event $s_{j}^{(k)}$ was considered a match to a ground-truth event $t_{i}^{(l)}$ if the absolute time difference was less than or equal to the predefined matching window $\Delta$, which was typically set to 1 millisecond:
$n^{\text{match}}_{l,k}=\#\{i:|t_{i}^{(l)}-s_{j}^{(k)}|\leq\Delta\text{ for some }j\}.$ (12)
2. Missed Events and False Positives: Using the number of matching events, we computed the number of missed events ($n^{\text{miss}}_{l,k}$) and false positives ($n^{\text{fp}}_{l,k}$) as follows:
$n^{\text{miss}}_{l,k}=N_{l}-n^{\text{match}}_{l,k},\quad n^{\text{fp}}_{l,k}=M_{k}-n^{\text{match}}_{l,k},$ (13)
where $N_{l}$ was the total number of ground-truth events in $l$, and $M_{k}$ was the total number of sorted events in $k$.
3. Accuracy: Accuracy ($a_{l,k}$) balanced the contributions of both missed events and false positives and was defined as:
$a_{l,k}=\frac{n^{\text{match}}_{l,k}}{n^{\text{match}}_{l,k}+n^{\text{miss}}_{l,k}+n^{\text{fp}}_{l,k}}.$ (14)
4. Precision and Recall: The precision ($p_{l}$) and recall ($r_{l}$) were calculated based on the best matching sorted unit $\hat{k}_{l}$ for each ground-truth unit $l$. The best match was determined as the sorted unit $k$ with the highest accuracy:
$\hat{k}_{l}=\arg\max_{k}a_{l,k}.$ (15)
To identify the best matching sorted unit $\hat{k}_{l}$ for each ground-truth unit $l$, the Hungarian Algorithm was applied to establish the best one-to-one correspondence between sorted and ground-truth units. Using this best match, precision and recall were defined as:
$p_{l}=\frac{n^{\text{match}}_{l,\hat{k}_{l}}}{n^{\text{match}}_{l,\hat{k}_{l}}+n^{\text{fp}}_{l,\hat{k}_{l}}},\quad r_{l}=\frac{n^{\text{match}}_{l,\hat{k}_{l}}}{n^{\text{match}}_{l,\hat{k}_{l}}+n^{\text{miss}}_{l,\hat{k}_{l}}}.$ (16)
A.6 Hyperparameter Search
We performed a grid search to optimize the hyperparameters for the SimSort detection and identification models. The hyperparameter ranges included the number of layers, hidden size, learning rate, number of attention heads, dropout rate, and augmentation settings. The final selected values are presented in Tables S2 and S3.
A.7 Infrastructures
Our simulation dataset are generated using CPU clusters for approximately 5000 CPU hours, which can be parallelized. Our models are trained on NVIDIA A100/V100 GPU, approximately 20 GPU hours for each time.
Table S2: Hyperparameters of SimSort detection model.
Hyperparameters Description Value
Num Layer Number of Transformer encoder layers 4
Input Size Dimensions of the input data 4
Batch Size Number of items processed in a single operation 128
Learning Rate Learning rate of the model 5e-4
Hidden Size Dimension of the Transformer encoder 256
nhead Number of Transformer attention heads 4
Dropout Dropout probability in Transformer encoder 0.2
Weight Decay Regularization parameter used to prevent overfitting 1e-5
Warmup Steps Steps to warm up the learning rate 4000
Num Epochs Number of epochs to run 20
Sigmoid Threshold Probability threshold for classifying spike events 0.97
Noise Level The level of noise applied to the input data 0.9
Maximum Time Jitter Maximum variation in timing (in terms of upsampled sampling points) 5
Amplitude Scale Range The range for scaling the amplitude of the input data [0, 1]
Transform Probability The probability to perform augmentations to input data 0.8
Max Channels The maximum number of data channels augmentations are applied to 4
Table S3: Hyperparameters of SimSort identification model.
Hyperparameters Description Value
Num Layer Number of GRU layers 1
Input Size Dimensions of the input data 4
Batch Size Number of items processed in a single operation 128
Learning Rate Learning rate of the model 1e-4
Hidden Size Latent dimention size of GRU 512
Weight Decay Regularization parameter used to prevent overfitting 1e-5
Num Epochs Number of epochs to run 10
$k$ components Number of components to reconstruct waveform using SVD 5
Noise Level The level of noise applied to the input data 4
Maximum Time Jitter Maximum variation in timing (in terms of upsampled sampling points) 50
Amplitude Scale Range The range for scaling the amplitude of the input data [0.5, 1]
Transform Probability The probability to perform augmentations to input data 0.5
Max Channels The maximum number of data channels augmentations are applied to 4
Appendix B Supplementary Figures
Figure S1: (a) Accuracy of spike sorting on the Hybrid dataset. SimSort significantly outperforms HerdingSpikes2 ($p=0.0000^{*}$), JRClust ($p=0.0000^{*}$), KiloSort2 ($p=0.0000^{*}$), and Tridesclous ($p=0.0001^{*}$) in static recordings, and also shows a significant advantage over IronClust ($p=0.0296^{*}$). There is no significant difference compared to KiloSort ($p=0.5023$), MountainSort4 ($p=0.1799$), and SpykingCircus ($p=0.1356$). In the drift condition, SimSort significantly outperforms HerdingSpikes2 ($p=0.0000^{*}$), JRClust ($p=0.0000^{*}$), KiloSort ($p=0.0285^{*}$), KiloSort2 ($p=0.0000^{*}$), MountainSort4 ($p=0.0004^{*}$), SpykingCircus ($p=0.0111^{*}$), and Tridesclous ($p=0.0007^{*}$), with no significant difference compared to IronClust ($p=0.3398$). (b) Accuracy of spike sorting on the WaveClus dataset. In the easy subset, SimSort shows no significant improvement compared to KiloSort ($p=0.0770$), KiloSort2 ($p=0.1569$), MountainSort4 ($p=0.6137$), and MountainSort5 ($p=0.4424$). In the difficult subset, SimSort significantly outperforms KiloSort ($p=0.0007^{*}$), KiloSort2 ($p=0.0004^{*}$), and MountainSort5 ($p=0.0449^{*}$), with no significant difference compared to MountainSort4 ($p=0.0935$). Statistical analysis was performed using a paired two-tailed $t$-test.
Figure S2: Performance comparison on Fifty test sets of IBL Neuropixels dataset.
Table S4: Performance comparison of spike identification on 50 test sets from IBL Neuropixels dataset. ARI values are presented as mean $\pm$ standard deviation. Statistical significance was assessed with paired t-tests: $p=0.0012$ for SimSort vs. CEED, $p=0.0554$ for SimSort vs. CEED+UMAP.
Methods Test sets (seed 0-49)
CEED 0.46 $\pm$ 0.09
CEED+UMAP 0.47 $\pm$ 0.10
[…]
Figure S3: Visualized comparison results of three example test sets from IBL Neuropixels dataset. For each example, the first column presents GMM clustering results on UMAP-embedded SimSort representations, with predicted labels (top) and true labels (bottom). The second column displays GMM clustering on UMAP-embedded CEED representations. The third column illustrates GMM clustering directly on CEED representations (as in the CEED paper).
Figure S4: Impact of dimensionality reduction and clustering algorithms on the hybrid dataset. MS denotes Mean Shift, SC denotes Spectral Clustering. Comparing different combinations of dimensionality reduction methods (None, PCA, $t$-SNE, UMAP) and clustering algorithms (GMM, KMeans, Mean Shift, and Spectral Clustering) on static (left) and drift (right) subset. Here, “None” indicates that raw features learned by SimSort were used directly without dimensionality reduction.
Appendix C More Showcases
In this section, we present additional visualizations from various datasets to provide a more comprehensive evaluation of SimSort’s performance. These visualizations include examples of spike detection, identification, and sorting tasks. The supplementary showcases illustrate key aspects such as the alignment between detected and ground-truth spikes and the clustering of spike waveforms.
Figure S5: Visualization of spike detection performance on an example recording in Hybrid dataset.
Figure S6: Zoomed-in visualization of spike detection performance on an example recording in Hybrid dataset.
Figure S7: Visualization of spike detection performance on an example recording in WaveClus dataset.
Figure S8: Zoomed-in visualization of spike detection performance on an example recording in WaveClus dataset.
Figure S9: Visualization of spike detection performance on an example recording in BBP L6 dataset.
Figure S10: Zoomed-in visualization of spike detection performance on an example recording in BBP L6 dataset.
Figure S11: Visualization of spike identification performance on hybrid static subset.
Figure S12: Visualization of spike identification performance on hybrid drift subset.
Figure S13: Visualization of spike sorting performance on hybrid static subset.
Figure S14: Visualization of spike sorting performance on hybrid drift subset.
Figure S15: Visualization of spike sorting performance on WaveClus dataset.
Figure S16: Visualization of spike sorting performance on BBP L6 dataset.
Figure S17: Visualization of spike sorting performance on real recording data.